OpenStudy (baldymcgee6):

Can someone explain integration by partial fractions, or refer me to a good website that will help me?

OpenStudy (anonymous):

Essentially, when you're given a rational expression in the integrand, you can split it up into several fractions, each one with a denominator that's a different factor of the original denominator.

OpenStudy (baldymcgee6):

hmmm

OpenStudy (baldymcgee6):

it is always for quotients?

OpenStudy (anonymous):

For instance,\[\frac{1}{x(x^2+1)} = \frac{A}{x} + \frac{B}{x^2+1.}\]With this, you can cross-multiply, set appropriate values for x to get values of A and B, and then replace the two new fractions where the old, ugly one was.

OpenStudy (anonymous):

I don't think it's always for quotients, but that's where the vast majority of them lie.

OpenStudy (baldymcgee6):

it is to just break of the original?

OpenStudy (baldymcgee6):

break up*

OpenStudy (anonymous):

I believe so, yep. I haven't seen one in an introductory calculus course that isn't for that purpose.

OpenStudy (baldymcgee6):

what education do you have?

OpenStudy (baldymcgee6):

if you dont mind me asking.. :)

OpenStudy (anonymous):

I'm going through multivariate calculus now, and I've studied about 3 or 4 single-variable calculus textbooks made for the Calc 1 and 2 levels in university... Or were you asking me where I go to school? :P

OpenStudy (baldymcgee6):

haha no, i was just asking what level of education

OpenStudy (anonymous):

Ah, gotcha. I'm entering my freshman year of college.

OpenStudy (baldymcgee6):

nice, good work.